Solve by the elimination method.
9x-9y=18
7y-3x=-8
-3x+7y=-8
9x-9y=18
-3x+7y=-8
-9x+21y=-24
9x-9y=18
12y=-6
y=???
I really need help with this one. Thank you so much!!!

Answer by rfer(16322)

y=-1/2
x=1.5
take one of the equations and change so you solve for one unknow in terms of the other
9x=9y+18
x=(9y+18)/9
x=y+2
----------------
7y-3(y+2)=-8
7y-3y-6=-8
4y=-2
y=-1/2
--------------
9x=9(-1/2)+18
9x=-4.5+18
x=13.5/9
x=1.5
----------------
7(-1/2)-3(1.5)=-8
-3.5-4.5=-8
I will let you do more proof if you want